OpenStudy (anonymous):

What is the value of the determinant? -3 1 -6 5

OpenStudy (anonymous):

-15+6

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

Thank you very much!

OpenStudy (salina):

[(-3*5)-(-6*1)]=-15+6 = -9